If a 25.0 mL solution of acetic acid of unknown concentration required 30.0 mL of 0.100M NaOH solution to be neutralized, what is the concentration of the unknown substance.
Answer #1

Balanced chemical equation is:

NaOH + CH3COOH ---> CH3COONa + H2O

Here:

M(NaOH)=0.1 M

V(NaOH)=30.0 mL

V(CH3COOH)=25.0 mL

According to balanced reaction:

1*number of mol of NaOH =1*number of mol of CH3COOH

1*M(NaOH)*V(NaOH) =1*M(CH3COOH)*V(CH3COOH)

1*0.1*30.0 = 1*M(CH3COOH)*25.0

M(CH3COOH) = 0.12 M

Answer: 0.120 M